# Desynchronization Index: A New Connectivity Approach for Exploring Epileptogenic Networks

### v2

@@ VI Conclusion
In this work, we developed a new computational framework for analyzing SEEG signals, with the final goal of identifying the EZ in patients with drug-resistant epilepsy. Our method, named DI, exploits the PTE model to estimate the effective connectivity between the cortical sites explored via SEEG and consider as most epileptogenic the SEEG channels that present an abrupt desynchronization in the period immediately preceding the seizure propagation. To evaluate its clinical utility, we implemented DI over a dataset of $11$ patients and compared its performance in identifying epileptogenic sites against the EI algorithm. Our results denoted that DI shows a higher sensitivity than EI in the detection of both the EZ and the PZ, and the best accuracy is obtained when combining the two algorithms as a unique detection tool.
@@
The proposed approach, integrated with other quantitative biomarkers, may constitute a key support for SEEG interpretation. Indeed, DI underscores signal modifications not visually evident, helping to identify those sites that contribute to the seizure propagation or inhibit the epileptic discharge during the inter-ictal period. In future work, we intend to clinically validate the DI algorithm over a wider dataset, possibly including data from different clinical research centers, and evaluate the potential of our computational framework for analyzing SEEG signals during the inter-ictal period.

### v3

@@ 6 Conclusion
In this work, we developed a new computational framework for analyzing SEEG signals, with the final goal of identifying EZ in patients with drug-resistant epilepsy. Our method, named DI, uses PTE to estimate the effective connectivity between the SEEG cortical sites and considers as epileptogenic the SEEG channels that present an abrupt desynchronization in the period immediately preceding the propagation of seizures. To evaluate its clinical utility, we implemented DI on a dataset of $11$ patients and compared its performance in identifying epileptogenic sites against the EI algorithm. The results showed that DI leads to a higher AUC than EI in the EZ identification ($AUC=0.80$ vs $AUC=0.74$), and the best performance ($AUC=0.87$) is obtained when combining the two algorithms as a unique detection tool.
@@
Our approach, integrated with other quantitative biomarkers, may constitute a key support for SEEG interpretation. The DI algorithm underscores signal modifications that are not visually evident, helping to identify those sites that contribute to the propagation of seizures or inhibit the epileptic discharge during the inter-ictal period. In future work, we intend to clinically validate the DI algorithm on a wider dataset, possibly including data from different clinical research centers, and evaluate the potential of our computational framework for analyzing SEEG signals during the inter-ictal period.